# Additional dimensions of space and time in the region of deep inelastic processes

### v1

@@ Abstract
We prove that the well-known Heisenberg uncertainty relations and Landau-Peierls uncertainty relations implicitly contain “hidden” angular variables, which belong to new uncertainty relations. Based on the obtained relations, we derive a formula for estimating the speed $U^{*}$ of a virtual particle in indirect measurements. We applied the theory of indirect measurements and the derived formula to estimate the module of the group velocity of virtual photons from the DIS HERA data. The HERA data indicate that the speed of virtual photons exceeds the speed of light $c$ in free space, $U^{*}>c$. The properties of virtual photons and a hypothetical tachyon particle are almost identical. It is found that in the realm of particle interaction, the new angular parameters are closely related to the type of the phase-space geometry and dimensionality of the space-time continuum. It is suggested that the problem of the normalization condition $U^{*}=c$ at $Q^{2}=0\,\rm{GeV}^{2}$ can be solved naturally within the framework of “Two-Time Physics” developed by I. Bars. 2T-physics is the theory with local symplectic $\mathrm{Sp(2,R)}$ gauge symmetry in phase-space and the space-time geometry of signature $\mathrm{(1+1^{\prime},d+1^{\prime})}$ with one extra time-like and one extra space-like dimensions.
@@ I Introduction
The goal of the present report is twofold. First, we derive a formula for estimating a speed of virtual particles. Second, we combine the theory of indirect measurements, the hardware resolution of the ZEUS detector [8, 9] and the obtained formula into a mathematical tool for evaluating the speed of virtual particles from experimental data. As an application, we present preliminary results on the speed of virtual photons ($\gamma^{*}$) using a small set of deep inelastic scattering (DIS) data from the HERA collider [10].
@@
Let us recall three phenomena: 1) The formation of a Mach cone when the speed of a body in a medium is close to or greater than the speed of sound in the medium; 2) The Cherenkov-Vavilov effect [11, 12, 13]; 3) In 1904-1905 A. Sommerfeld [14] established in the context of Lorentz’s theory of electromagnetism [15] that if the speed of the electron $\mathrm{U}$ is less than the speed of light in a vacuum, $\mathrm{U}<c$, then the electron is able to move at a constant velocity. However, if $\mathrm{U}>c$, then an external force is required for uniform motion.
@@
In the next two paragraphs we present derivations of new uncertainty relations and a formula for evaluating the speed of a virtual particle in indirect measurements. After presenting experimental results for the lower bound on the speed of virtual photons, we discuss the tachyonic properties of $\gamma^{*}$, followed by arguments that allow relating the superluminal speed of virtual photons to the metric properties of space-time (in particular, its dimensionality).
@@ III Estimation of a particle speed from the uncertainty relations
Here $A_{t}=3{\delta^{2}\!\!\!}_{{}_{LP}}/({\delta^{2}\!\!\!}_{{}_{H}}\,\cos{\psi}_{{}_{H}})$ is the theoretical magnitude of the normalization parameter. In the next paragraph, we will see how the velocity of virtual particles relates the value of $A_{t}$ and the metric properties of space-time.
IV Space-time metrics in the interaction region
@@
Figure 1: (a) Kinematic ($x_{{}_{Bj}}$, $Q^{2}$) grid of the combined HERA I data for neutral current $e^{+}p$ deep inelastic scattering events at the center-of-mass energy $\sqrt{s_{com}}\approx 300$ GeV [10], Table 11. Different colors indicate data with the inelasticity, $y=Q^{2}/x_{{}_{Bj}}{s}_{com}$, in the following intervals : $0.354-0.47$ (red); $0.47-0.511$ (blue); $0.511-0.598$ (black); $0.598-0.676$ (magenta); $0.676-0.951$ (green). (b) The normalized virtual photon speed, $\beta^{*}=\|\mathrm{U}_{lb}^{*}\|/c$, as a function of the photon virtuality $Q^{2}$. The color of the dots corresponds to different values of the inelasticity $y$ as in Fig (a). It follows from the condition, $\beta^{*}(Q^{2}\rightarrow 0)=$1 that the normalization parameter $A_{e}=1/\sqrt{3}$. (c) The same as in panel (b), but for $\beta^{*}$ as a function of two variables $x_{{}_{Bj}}$ and $Q^{2}$ at different $y$-intervals. (d) The same as in panel (c), but the figure is rotated so that the data points are roughly projected onto a curve.
Let us first discuss possible values of the parameter $A_{t}$ in formula (9), and denote by $A_{e}$ its value found from the data. The Standard Model assumes that the geometry of space-time known from macroscopic physics also holds in the microcosm too. The derivation of the inequality (4) was based on this assumption, using the dot product of vectors in proper Euclidean space. The inequality (6) was derived from the inequality (5). Therefore, there is a good reason to believe that ${\delta^{2}\!\!\!}_{{}_{LP}}={\delta^{2}\!\!\!}_{{}_{H}}$. In this case, $A_{t}=3/\cos{\psi}_{{}_{H}}\geq 3$. Consequently, if it follows from an experimental data that $A_{e}\geq 3$, then the interaction of particles takes place in the region of space-time with the Minkowski geometry and the space index of $k=3$.
The case is quite different if $A_{e}<3$. Then our assumption about the metric of space-time in the interaction region is not correct.
@@
As input, we use the combined data from the H1 and ZEUS experiments on deep inelastic $ep$ scattering at the HERA collider [10]. The magnitude of the photon virtuality, $Q^{2}$, varies over a very wide range of values. Neutral current interaction cross sections at low $Q^{2}\leq$100 GeV2 are dominated by virtual photon exchange. In the limit $Q^{2}\rightarrow 0$ GeV2 (a real photon limit), $\beta^{*}=\|\mathrm{U}_{lb}^{*}/c\|\rightarrow 1$ should hold. This condition allows to fix the value of $A_{e}$ and at $Q^{2}>0$ GeV2 to set a lower bound on the speed of virtual photons.
@@
Figure 1(b) shows the speed of virtual photons normalized to the speed of real photons, $\beta^{*}=\|\mathrm{U}_{lb}^{*}\|/c$, as a function of $Q^{2}$ at different $y$-intervals. The result follows from Eqs. (9) - (13) and the HERA I data for neutral current $e^{+}p$ deep inelastic scattering events with beam momenta $(l,P)=$(27.5, 820) GeV/c and the center-of-mass energy $\sqrt{s_{com}}\approx 300$ GeV [10], Table 11. Note that $\beta^{*}$ grows with $Q^{2}$, but the slope of this growth decreases with $y$. The condition $\beta^{*}(Q^{2}\rightarrow 0)=$1 fixes $A_{e}$. In this way we get $\sqrt{3}A_{e}=1$. Figures 1(c) and 1(d) show the dependence of $\beta^{*}$ on two variables, $x_{{}_{Bj}}$ and $Q^{2}$. And again we see that $\beta^{*}$ grows with $Q^{2}$ and $x_{{}_{Bj}}$. Figure 1(d) is rotated view of Fig. 1(c) in order to project all data points on the same curve. This demonstrates that almost all points at different $y$ are located on a flat surface.
The results presented in Fig. 1 show that the speed of virtual photons at $Q^{2}>0$ GeV2 exceeds the speed of light $c$ in free space, $\beta^{*}>1$. Superluminal virtual photons can be interpreted as representatives of a class of superluminal particles, the tachyons [24, 25, 26, 27, 28, 29], since their properties are largely similar as shown in Table 1,
Table 1: Comparison of the properties of tachions and virtual photons.
Tachyon Virtual photon, $\gamma^{*}$
mass $(m^{*})^{2}<0$ $(m^{*})^{2}=-Q^{2}<0$
energy $\epsilon<0,>0$ $q_{0}<0,>0$
@@
$\epsilon\rightarrow 0,\beta^{*}\rightarrow\infty$ $q_{0}\rightarrow 0,\beta^{*}\rightarrow\infty$ (?)
@@
I) Nonlinear fields in 4D [30, 31]. In nonlinear field theories (with scalar or/and vector fields) the signal velocity depends on the magnitude of the field and its derivatives. In this case, the speed of the signals can be less or more than the speed of light $c$ in void and the space index change from $k=3$ to $k=2$ or even to $k=1$ and hence, $A_{t}<3$. This corresponds to the space-time domain, in which there are two or three time-like dimensions. However, in this scenario there are problems not only with causality (paradoxes), the violation of unitarity and occurrence of ghost modes with negative norm (for more discussion see Refs. [32, 33]), but also the entire relativistic kinematics in the Minkowski space is destroyed.
II) Stochastic time. One can increase the pseudo-Euclideanity of space-time by adding an additional time-like dimension in the interaction region, i.e., going from signature (1,3) to signature (2,3). Analysis of this scenario shows [33], that in the case of a space-time with a thermally excited second time dimension the dynamics of the physical system is diffusive, not ballistic, so that all trajectories involving motion in the second time dimension are dynamically unstable, thus allowing us to avoid the difficulties outlined for nonlinear field theories. The thermal extra time dimension behaves like an extra spatial dimension and the high temperature limit for the thermal extra time dimension is equivalent to a small compactification radius for an extra spatial dimension. Therefore, such a scenario is consistent with Kaluza-Klein type theories. However, simply adding another spatial or temporal dimension does not solve the problem we encountered ($A_{t}>3$, $A_{e}<3$).
@@
The principle of local gauge invariance is an important ingredient in the construction of realistic models for the interactions observed in nature. I. Bars with co-authors [38], [39] [32] discovered a fundamental role of the local symplectic Sp(2, R) gauge symmetry in phase space, which gave rise to new field theories in (2,4) space-time with one extra time and one extra space dimensions (2T-physics). This includes 2T field theories that yield one-time (1T) field theories for the Standard Model [40], General Relativity [41], SUSY [42] and others. The canonical transformations of the type Sp(2, R), considered as a local gauge symmetry, has the power to cure the ghost and causality problems of extra time-like dimensions.
In the following, we give a qualitative argument for the fact that the condition $A_{t}<3$ is feasible in the framework of 2T-physics. First, we postulate that the interaction region of DIS has a small spatial and temporal extent, and within the framework of the 2T-physics, the interactions take place in $\mathrm{(1+1^{\prime},3+1^{\prime})}$ space with one extra time and one extra space dimensions and the gauge invariant sector of 2T-physics, namely the ghost free physical sector, effectively becomes a 1T theory with an effective 1+3 dimensions [43]. Mathematically, the speed of a photon in the physical sector is a combination of the three effective spatial dimensions and one effective temporal dimension (the ”effective Minkowski space-time”). So, we assume that during the DIS process a virtual photon “captures” (or “admixes”) “new” additional spatial $x$ or time-like $\tau$ dimensions with the probability $\alpha$ and captures the “old” space-time dimensions, $\{t,x^{i}\}$, with the probability $\omega$. We label the components containing the conventional time dimension with index $H$, and the components containing an additional time-like dimension with index $B$. As a result, the squares of the normalized photon velocities as spatio-temporal combinations give contributions with the following probabilities: ${\beta^{2}}_{H}(t,x^{1},x^{2},x^{3})\sim\omega$, $\tilde{\beta^{2}}_{H}(t,x^{i},x^{j},x)\sim 3\alpha$, ${u^{2}}_{B}(\tau,x^{1},x^{2},x^{3})\sim\alpha$ and $\tilde{u^{2}}_{B}(\tau,x^{i},x^{j},x)\sim 3\alpha^{2}$. The total probability must satisfy the condition:
@@
In the limit $Q^{2}\to 0$, all speeds in Eq. (15) should be equated to unity. As a result, we get the normalization condition $\sqrt{3}A_{eff}=1$,
@@
These restrictions on $\cos{\psi}_{{}_{H}}$ and $\cos{\psi}_{{}_{B}}$ are quite soft. Thus, in the space-time volume of a high-energy reaction a “mixing in” of extra dimensions is possible and the normalization condition $\sqrt{3}A_{eff}=1$ is feasible. The specific values of $\cos{\psi}_{{}_{H}}$, $\cos{\psi}_{{}_{B}}$, $\alpha$, ${\beta^{2}}_{H}$, $\tilde{\beta^{2}}_{H}$, $\tilde{u^{2}}_{B}$, $\tilde{u^{2}}_{B}$ can be found by fitting Eq. (15) to the data similar to Fig. 1(c).

### v2

# Additional dimensions of space and time in the domain of deep inelastic processes

@@ Abstract
We prove that the well-known Heisenberg uncertainty relations and Landau-Peierls uncertainty relations implicitly contain “hidden” angular variables, which belong to new uncertainty relations. Based on the obtained relations, we derive a formula for estimating speed $U^{*}$ of a virtual particle in indirect measurements. We applied the theory of indirect measurements and the derived formula to estimate the module of the group velocity of virtual photons from the DIS HERA data. The HERA data indicate that the speed of virtual photons exceeds the speed of light $c$ in free space, $U^{*}>c$. The properties of virtual photons and a hypothetical tachyon particle are almost identical. It is found that in the realm of particle interaction, the new angular parameters are closely related to the type of the phase-space geometry and dimensionality of the spacetime continuum. It is suggested that the problem of the normalisation condition $U^{*}=c$ at $Q^{2}=0\,\rm{GeV}^{2}$ can be solved naturally within the framework of “Two-Time Physics” developed by I. Bars. 2T-physics is the theory with local symplectic $\mathrm{Sp(2,R)}$ gauge symmetry in phase-space and the spacetime geometry of signature $\mathrm{(1+1^{\prime},d+1^{\prime})}$ with one extra time-like and one extra space-like dimensions.
@@ I Introduction
The goal of the present communication is twofold. First, we derive a formula for estimating a speed of virtual particles. Second, we combine the theory of indirect measurements, the hardware resolution of the ZEUS detector [8, 9] and the obtained formula into a mathematical tool for evaluating the speed of virtual particles from experimental data. As an application, we present preliminary results on the speed of virtual photons ($\gamma^{*}$) using a small set of deep inelastic scattering (DIS) data from the HERA collider [10].
@@
Let us recall three phenomena: 1) The formation of a Mach cone when the speed of a body in a medium is close to or greater than the speed of sound in the medium; 2) The Cherenkov-Vavilov effect [11, 12, 13]; 3) In 1904-1905 A. Sommerfeld [14] established in the context of Lorentz’s theory of electromagnetism [15] that if the speed of the electron $\mathrm{U}$ is less than the speed of light in a vacuum, $\mathrm{U}<c$, then the electron is able to move at a constant velocity. However, if $\mathrm{U}>c$, then an external force is required for uniform motion. It is not superfluous to note that most of these conclusions were made much earlier by O. Heaviside [16, 17].
@@
In the next two sections we present derivations of new uncertainty relations and a formula for evaluating the speed of a virtual particle in indirect measurements. After presenting experimental results for the lower bound on the speed of virtual photons, we discuss the tachyonic properties of $\gamma^{*}$, followed by arguments that allow relating the superluminal speed of virtual photons to the metric properties of spacetime (in particular, its dimensionality).
@@ III Estimation of a particle speed from the uncertainty relations
Here $A_{t}=3{\delta^{2}\!\!\!}_{{}_{LP}}/({\delta^{2}\!\!\!}_{{}_{H}}\,\cos{\psi}_{{}_{H}})$ is the theoretical magnitude of the normalisation parameter. In the next section, we will see how the velocity of virtual particles relates the value of $A_{t}$ and the metric properties of spacetime.
IV Spacetime metrics in the interaction domain
@@
Figure 1: (a) Kinematic ($x_{{}_{Bj}}$, $Q^{2}$) grid of the combined HERA I data for neutral current $e^{+}p$ deep inelastic scattering events at the center-of-mass energy $\sqrt{s_{com}}\approx 300$ GeV [10], Table 11. Different colors indicate data with the inelasticity, $y=Q^{2}/x_{{}_{Bj}}{s}_{com}$, in the following intervals : $0.354-0.47$ (red); $0.47-0.511$ (blue); $0.511-0.598$ (black); $0.598-0.676$ (magenta); $0.676-0.951$ (green). (b) The normalized virtual photon speed, $\beta^{*}=\|\mathrm{U}_{lb}^{*}\|/c$, as a function of the photon virtuality $Q^{2}$. The color of the dots corresponds to different values of the inelasticity $y$ as in Fig (a). It follows from the condition, $\beta^{*}(Q^{2}\rightarrow 0)=$1 that the normalisation parameter $A_{e}=1/\sqrt{3}$. (c) The same as in panel (b), but for $\beta^{*}$ as a function of two variables $x_{{}_{Bj}}$ and $Q^{2}$ at different $y$-intervals. (d) The same as in panel (c), but the figure is rotated so that the data points are roughly projected onto a curve.
Let us first discuss possible values of the parameter $A_{t}$ in formula (9), and denote by $A_{e}$ its value found from the data. The Standard Model assumes that the geometry of spacetime known from macroscopic physics also holds in the microcosm too. The derivation of the inequality (4) was based on this assumption, using the dot product of vectors in proper Euclidean space. The inequality (6) was derived from the inequality (5). Therefore, there is a good reason to believe that ${\delta^{2}\!\!\!}_{{}_{LP}}={\delta^{2}\!\!\!}_{{}_{H}}$. In this case, $A_{t}=3/\cos{\psi}_{{}_{H}}\geq 3$. Consequently, if it follows from an experimental data that $A_{e}\geq 3$, then the interaction of particles takes place in the domain of spacetime with the Minkowski geometry and the space index of $k=3$.
The case is quite different if $A_{e}<3$. Then our assumption about the metric of spacetime in the interaction domain is not correct.
@@
As input, we use the combined data from the H1 and ZEUS experiments on deep inelastic $ep$ scattering at the HERA collider [10]. The magnitude of the exchange particles virtuality, $Q^{2}$, varies over a very wide range of values. Neutral current interaction cross sections at low $Q^{2}\leq$100 GeV2 are dominated by the virtual photon exchange. In the limit $Q^{2}\rightarrow 0$ GeV2 (a real photon limit), $\beta^{*}=\|\mathrm{U}_{lb}^{*}/c\|\rightarrow 1$ should hold. This condition allows to fix the value of $A_{e}$ and at $Q^{2}>0$ GeV2 to set a lower bound on the speed of virtual photons.
@@
Figure 1(b) shows the speed of virtual photons normalized to the speed of real photons, $\beta^{*}=\|\mathrm{U}_{lb}^{*}\|/c$, as a function of $Q^{2}$ at different $y$-intervals. The result follows from Eqs. (9) - (13) and the HERA I data for neutral current $e^{+}p$ deep inelastic scattering events with beam momenta $(l,P)=$(27.5, 820) GeV/c and the center-of-mass energy $\sqrt{s_{com}}\approx 300$ GeV [10], Table 11. Note that $\beta^{*}$ grows with $Q^{2}$, but the slope of this growth decreases with $y$. The condition $\beta^{*}(Q^{2}\rightarrow 0)=$1 fixes $A_{e}$. In this way one get $\sqrt{3}A_{e}=1$. Figures 1(c) and 1(d) show the dependence of $\beta^{*}$ on two variables, $x_{{}_{Bj}}$ and $Q^{2}$. And again we see that $\beta^{*}$ grows with $Q^{2}$ and $x_{{}_{Bj}}$. Figure 1(d) is rotated view of Fig. 1(c) in order to project all data points on the same curve. This demonstrates that almost all points at different $y$ are located on a flat surface.
The results presented in Fig. 1 show that the speed of virtual photons at $Q^{2}>0$ GeV2 exceeds the speed of light $c$ in free space, $\beta^{*}>1$. A perusal of the litrature on faster-than-light particles reveals that virtual photons can be interpreted as representatives of a class of superluminal particles, hypothetical tachyons [26, 27, 28, 29, 30, 31], since their properties are largely similar as shown in Table 1,
Table 1: Comparison of the properties of virtual photons and tachyons.
Virtual photon, $\gamma^{*}$ Tachyon
mass $(m^{*})^{2}=-Q^{2}<0$ $(m^{*})^{2}<0$
energy $q_{0}<0,>0$ $\epsilon<0,>0$
@@
$q_{0}\rightarrow 0,\beta^{*}\rightarrow\infty$ (?) $\epsilon\rightarrow 0,\beta^{*}\rightarrow\infty$
Moreover, such a result has been expected for a long time. Thanks to many years of extensive theoretical studies, mainly carried out by a group of Italian physicists led by E. Recami (for an overview see [32, 33]), it became clear that tachyons (or spacelike states) must exist as intermediate states or exchanged objects in elementary processes [34, 35, 32]. For reviews of searches for evidence of tachyons see [36, 37].
Thus, using a different theoretical approach and based on modern experimental data from high-energy physics, we have confirmed the predictions made more than 40 years ago.
@@
I) Nonlinear fields in 4D [38, 39]. In nonlinear field theories (with scalar or/and vector fields) the signal velocity depends on the magnitude of the field and its derivatives. In this case, the speed of the signals can be less or more than the speed of light $c$ in void and the space index change from $k=3$ to $k=2$ or even to $k=1$ and hence, $A_{t}<3$. This corresponds to the spacetime domain, in which there are two or three time-like dimensions. However, in this scenario there are problems not only with causality (paradoxes), the violation of unitarity and occurrence of ghost modes with negative norm (for more discussion see Refs. [40, 41]), but also the entire relativistic kinematics in the Minkowski space is destroyed.
II) Stochastic time. One can increase the pseudo-Euclideanity of spacetime by adding an additional time-like dimension in the interaction domain, i.e., going from signature (1,3) to signature (2,3). Analysis of this scenario shows [41], that in the case of a spacetime with a thermally excited second time dimension the dynamics of the physical system is diffusive, not ballistic, so that all trajectories involving motion in the second time dimension are dynamically unstable, thus allowing us to avoid the difficulties outlined for nonlinear field theories. The thermal extra time dimension behaves like an extra spatial dimension and the high temperature limit for the thermal extra time dimension is equivalent to a small compactification radius for an extra spatial dimension. Therefore, such a scenario is consistent with Kaluza-Klein type theories. However, simply adding another spatial or temporal dimension does not solve the problem we encountered ($A_{t}>3$, $A_{e}<3$).
@@
The principle of local gauge invariance is an important ingredient in the construction of realistic models for the interactions observed in nature. I. Bars with co-authors [46], [47], [40] discovered a fundamental role of the local symplectic Sp(2, R) gauge symmetry in phase space, which gave rise to new field theories in (2,4) spacetime with one extra time and one extra space dimensions (2T-physics). This includes 2T field theories that yield one-time (1T) field theories for the Standard Model [48], General Relativity [49], SUSY [50] and others. The canonical transformations of the type Sp(2, R), considered as a local gauge symmetry, has the power to cure the ghost and causality problems of extra time-like dimensions.
In the following, we give a qualitative argument for the fact that the condition $A_{t}<3$ is feasible in the framework of 2T-physics. First, we postulate that the interaction domain of DIS has a small spatial and temporal extent, and within the framework of the 2T-physics, the interactions take place in $\mathrm{(1+1^{\prime},3+1^{\prime})}$ space with one extra time and one extra space dimensions and the gauge invariant sector of 2T-physics, namely the ghost free physical sector, effectively becomes a 1T theory with an effective 1+3 dimensions [51]. Mathematically, the speed of a photon in the physical sector is a combination of the three effective spatial dimensions and one effective temporal dimension (the ”effective Minkowski spacetime”). So, we assume that during the DIS process a virtual photon “captures” (or “perceives”) the “new” additional spatial $x$ or time-like $\tau$ dimensions with the probability $\alpha$ and captures the “old” spacetime dimensions, $\{t,x^{i}\}$, with the probability $\omega$. We label the components containing the conventional time dimension with index $H$, and the components containing an additional time-like dimension with index $B$. As a result, the squares of the normalized photon velocities as spatio-temporal combinations give contributions with the following probabilities: ${\beta^{2}}_{H}(t,x^{1},x^{2},x^{3})\sim\omega$, $\tilde{\beta^{2}}_{H}(t,x^{i},x^{j},x)\sim 3\alpha$, ${u^{2}}_{B}(\tau,x^{1},x^{2},x^{3})\sim\alpha$ and $\tilde{u^{2}}_{B}(\tau,x^{i},x^{j},x)\sim 3\alpha^{2}$. The total probability must satisfy the condition:
@@
In the limit $Q^{2}\to 0$, all speeds in Eq. (15) should be equated to unity. As a result, we get the normalisation condition $\sqrt{3}A_{eff}=1$,
@@
These restrictions on $\cos{\psi}_{{}_{H}}$ and $\cos{\psi}_{{}_{B}}$ are quite soft. Thus, in the spacetime volume of a high-energy reaction a “mixing in” of extra dimensions is possible and the normalisation condition $\sqrt{3}A_{eff}=1$ is feasible. The specific values of $\cos{\psi}_{{}_{H}}$, $\cos{\psi}_{{}_{B}}$, $\alpha$, ${\beta^{2}}_{H}$, $\tilde{\beta^{2}}_{H}$, $\tilde{u^{2}}_{B}$, $\tilde{u^{2}}_{B}$ can be found by fitting Eq. (15) to the data similar to Fig. 1(c).